ANOVA df SS Regression 1 0.72 Residual 10 62.6 Total 11 63.32 Coefficients Std Error Intercept...

ANOVA

df

SS

Regression

1

0.72

Residual

10

62.6

Total

11

63.32

Coefficients

Std Error

Intercept

14.64

146.76

No. of accounts (000)

1.99

5.87



This printout is for data relating the number of ATM withdrawals (in thousands) to the number of accounts (in thousands) at that branch. Predict the number of withdrawals if the number of accounts is 24.528 thousand. State the answer in thousands correct to two decimal places.
Answer #1

The regression model can be written as,

Y= 14.96+1.99(no.of accounts)

Then we can predict the no.of withdrawals at x=24.528

Y ( predicted)=14.96+1.99(24.528)

=14.96+40.81

Y( predicted)= 89.77=~90 ATM withdrawls